OpenStudy (anonymous):

HELP!!! Solve the equation. Check for extraneous solutions. |4x+3|=9+2x x= _________ or _________

OpenStudy (dls):

x=3

OpenStudy (dls):

and ummmm

OpenStudy (dls):

x=4

OpenStudy (anonymous):

Is that what i have to put in ans solve for? or is that the answers?

OpenStudy (dumbcow):

separate it into 2 equations when there is an absolute value \[4x+3 = \pm (9+2x)\] solve 4x+3 = 9+2x and 4x+3 = -(9+2x)

OpenStudy (dls):

i wrote the answer!

OpenStudy (dumbcow):

x=4 does not work

OpenStudy (anonymous):

How do i start to solve them?

OpenStudy (dls):

not sure about the 2nd value

OpenStudy (dumbcow):

just get x by itself by combining like terms example: 3x-1 = 5 add 1 to both sides to move 1 to right side divide by 3 to both sides to move 3 to right side

OpenStudy (anonymous):

What do i subtract first?

OpenStudy (anonymous):

4x+3=9+2x. -3 from both sides?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

and that leaves me with 4x=6+2x. What then?

OpenStudy (dumbcow):

combine the "x" terms

OpenStudy (dumbcow):

move 2x to left side

OpenStudy (anonymous):

So.. 6x=6 divide by 6?

OpenStudy (anonymous):

so... 1?

OpenStudy (dumbcow):

wait...you added , always do the inverse(opposite) operation

OpenStudy (anonymous):

How?

OpenStudy (dumbcow):

4x = 6 +2x --> subtract 2x 4x = 6-2x --> add 2x

OpenStudy (dumbcow):

same reason you took -3 from both sides earlier ....to cancel the 3 on right side now take "-2x" to cancel the 2x on right side

OpenStudy (anonymous):

2x=6? then x=3?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

is the other one x=-6

OpenStudy (dumbcow):

nope

OpenStudy (anonymous):

Well i dont know then. I tried

OpenStudy (dumbcow):

ok lets see where you messed up 4x+3 = -(9+2x) distribute 4x+3 = -9 -2x -3 to both sides 4x = -12 -2x 2x to both sides 6x = -12 divide x = -2

OpenStudy (anonymous):

Oh. i didnt put a negative on 2x. So my answers are x=3? and x=-2?

OpenStudy (dumbcow):

don't forget to distribute the negative yes those are the answers

OpenStudy (anonymous):

Thank you